show that J=oE is another representation of Ohm's law.
Answer #1

Let the current flowing through the conductor be I (ampere) when the applied potential difference is v (volt)

we know that from ohm's law

I=\frac{V}{R}

also,

R=\rho \frac{L}{A}

Then,

I=\frac{V}{\rho \frac{L}{A}}

\frac{I}{A}=\frac{V}{\rho L}   \left ( Current density (J)=\frac{I}{A} ; Electric field(E)=\frac{V}{L};Conductivity(\sigma)=\frac{1}{\rho }\right )

J=\sigma E